# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ 3.2. Global well-posedness
$\left\|\mathcal{N}_{2,q}(\psi_{1})-\mathcal{N}_{2,q}(\psi_{2})\right\|_{N^{0}}\leq\|\psi_{1}-\psi_{2}\|_{L_{t}^{1}L_{x}^{2}}+\|(|\psi_{1,q}|^{2\alpha}+|\psi_{2,q}|^{2\alpha})|c_{1}-c_{2}|\|_{L_{t}^{\frac{4}{3}}L_{x}^{\frac{3}{2}}}$
$+\|(|\psi_{1,q}|^{2\alpha}+|\psi_{2,q}|^{2\alpha})|v_{1}-v_{2}|\|_{L_{t}^{\frac{2}{3-\alpha}}L_{x}^{\frac{6}{2\alpha+1}}}$
@@
We now prove continuous dependence on the initial data. As in the proof of Proposition 3.2, we rely on a auxiliary metric to compensate for the lack of regularity of the nonlinearity $f$ and to deal with the non-integrability of the wave-functions. However, by virtue of Lemma 4.2, it suffices to consider the affine problem (4.3). This decomposition enables us to implement an argument in $L^{2}([0,T];L^{6}({\mathbf{R}}^{3}))$. In particular, it is sufficient to prove sequential continuity.
@@
where we used (2.25) in the second last inequality and (2.21) to control the difference of the free solutions in the last inequality. More precisely,
@@
To bound the second contribution in (4.5), we proceed as in (3.21). More precisely, we observe that (3.24) remains valid upon replacing the admissible Strichartz pair $(4,4)$ for $d=2$ by $(\frac{8}{3},4)$ for $d=3$. Hence, the respective version of (3.24) reads that there exists $\theta_{2}\in(0,1]$ such that
@@
Summing up and applying the Strichartz estimate (2.25), we conclude from Lemma 4.4 that there exists $C=C(M)>0$ and $\theta>0$ such that
@@
as $n\rightarrow\infty$ and where $(q,r)$ as in (4.1). The proof follows closely the one of (3.17) to which we refer for full details. In view of the Strichartz estimates of Lemma 2.14 it follows
@@
$\left|\nabla\mathcal{N}(\psi)\right|\leq C(1+|\psi|^{2\alpha})|\nabla\psi|\leq C(1+|\psi_{q}|^{2\alpha})|\nabla\psi|.$
More precisely, for $G_{\infty},G_{q}$ defined in (2.38) and upon applying (2.25), we split the non-homogeneous term in
@@
$\leq\|(G_{\infty})(\psi)|\nabla v-\nabla v_{n}|\|_{L_{t}^{1}L_{x}^{2}}+\|G_{q}(\psi)|\nabla v-\nabla v_{n}|\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}$
$+\left\|\left(G_{\infty}(\psi)-G_{\infty}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi)-G_{q}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}$
@@
$+\left\|\left(G_{\infty}(\psi)-G_{\infty}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi)-G_{q}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}.$
@@
$\|\nabla v-\nabla v_{n}\|_{L^{\infty}([0,T_{2}],L^{2}({\mathbf{R}}^{3}))}+\|\nabla v-\nabla v_{n}\|_{L^{q}([0,T_{2}],L^{r}({\mathbf{R}}^{3}))}\leq C\delta(c+v_{0},c_{n}+v_{0}^{n})\\ +\left\|\left(G_{\infty}(\psi)-G_{\infty}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi)-G_{q}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}.$
To conclude that (4.7) holds, it suffices to show that the second line of the right-hand side converges to 00 as $n$ goes to infinity. We proceed by contradiction assuming that there exists a subsequence still denoted $\psi_{n}$ such that there exists $\varepsilon>0$ such that for all $n$ sufficiently large,
(4.10) $\left\|\left(G_{\infty}(\psi)-G_{\infty}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi)-G_{q}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\geq\varepsilon.$
Inequality (4.5) implies that up to extracting a further subsequence, still denoted $\psi_{n}$, that $\psi_{n}=c_{n}+v_{n}$ converges to $\psi=c+v$ a.e. on $[0,T)\times{\mathbf{R}}^{3}$. By virtue of the Assumption 1.1, one has that $G_{\infty},G_{q}$ are continuous. Therefore,
[…]
@@
$\|G_{q}(\psi_{n})\|_{L_{t}^{\infty}L_{x}^{\frac{2(\alpha+1)}{2\alpha}}({\mathbf{R}}^{3})}\leq C\||\psi_{n}|^{2\alpha}(1-\chi(\psi_{n}))\|_{L_{t}^{\infty}L_{x}^{\frac{2(\alpha+1)}{2\alpha}}({\mathbf{R}}^{3})}$
@@
for all $n\in{\mathbf{N}}$, where we exploited (2.8), namely that the measure of $\operatorname{supp}(1-\chi(\psi_{n}))$ is finite. We obtain that there exists $\phi\in L^{\infty}([0,T];L^{2(\alpha+1)}({\mathbf{R}}^{3}))$ such that $|\psi_{q,n}|\leq\phi$ a.e. on $[0,T)\times{\mathbf{R}}^{3}$. Therefore, we control
$\left|\left(G_{\infty}(\psi)-G_{\infty}(\psi_{n})\right)\right|\,|\nabla v|\leq C|\nabla\psi|\in L^{1}([0,T);L^{2}({\mathbf{R}}^{3})),$
$\left|\left(G_{q}(\psi)-G_{q}(\psi_{n})\right)\right|\,|\nabla v|\leq C\left(|\psi|^{2\alpha}+|\phi|^{2\alpha}\right)|\nabla\psi|\in L^{q^{\prime}}([0,T);L^{r^{\prime}}({\mathbf{R}}^{3})).$
@@
Next, we show a persistence of regularity property and that the Hamiltonian energy $\mathcal{H}$ is conserved for regular solutions. Even though the proof is completely analogous to one for $d=2$, except that here we can exploit the affine structure of the energy space $\mathbb{E}$ and Sobolev embeddings depend on the dimension. For the sake of clarity, we provide the proof of this lemma.
@@
By means of the Strichartz estimates of Lemma 2.14, it follows for the admissible pair $(q,r)$ as in (4.1) and any $0<T<T^{\ast}$ that
@@
with $G_{1},G_{2}$ defined in (2.37). Upon splitting $G_{i}$ in $G_{i,\infty}$ and $G_{i,q}$, as in (2.38), it follows that
@@
where we note that $10<\alpha-1<$ 1. It follows
@@
Finally, we conclude that $\mathcal{H}(c+v)(t)=\mathcal{H}(c+v_{0})$ by performing the analogue argument as in the proof of Lemma 3.5 for $d=2$. ∎
@@
It only remains to show that the Hamiltonian energy is conserved for all solutions $\psi\in C([0,T^{\ast}),\mathbb{E}({\mathbf{R}}^{3}))$ which follows from Proposition 4.1, approximation by smooth solutions by means of Lemma 2.10 together with Lemma 4.5. ∎
@@ 4.1. Global well-posedness
Exploiting the affine structure of the energy space $\mathbb{E}({\mathbf{R}}^{3})$, we also prove global well-posedness for a class of equations for which the associated nonlinear potential energy density $F(|\psi|^{2})$ fails to be non-negative. Such equations arise for instance in nonlinear optics to investigate self-focusing phenomena in a defocusing medium, see [5, 46, 58]. A showcase model for such phenomena is (4.3) with competing subcritical power-type nonlinearities satisfying Assumption 1.5 and further of the form
@@
where $a_{1},a_{2}>0$ and $20<\alpha_{2}<\alpha_{1}<$ 2. The defocusing nonlinearity is dominant for large intensities $|\psi|^{2}>>\rho_{0}$ and focusing phenomena occur for small intensities $|\psi|^{2}\leq\rho_{0}$ where $\rho_{0}$ is determined by the far-field. The case $\alpha_{1}=2$, $\alpha_{2}=1$ corresponds to the energy-critical cubic-quintic nonlinearity and is investigated in [42, 44]. As before, we set $\rho_{0}=1$, further as in (4.4), it suffices to consider $c=1$ and upon scaling space and time $a_{1}=1$. We are hence led to consider nonlinearities of the type
@@
where Assumption 1.5 implies $\frac{\alpha_{1}}{\alpha_{2}}>a_{2}$. Furthermore, we may assume that $a_{2}>1$ as otherwise $F\geq 0$. Indeed,
@@
    if $a_{2}>1$, then $f$ admits two positive real roots $\rho_{1},1$ with $10<\rho_{1}<$ 1 and $F$ displays a local minimum in $\rho=1$ and a local maximum in $\rho=\rho_{1}$. Depending on the location of the root $\rho_{1}$ two scenarios may occur:
@@
      the root $\rho_{1}$ is sufficiently close to 00 such that $F(\rho)\geq 0$ for all $\rho\geq 0$,
@@
Let $f$ be a real-valued function satisfying Assumption 1.5 and further of the form
@@
with $20<\alpha_{1}<$ 2 and where $g\in C^{0}([0,\infty))\cap C^{1}(0,\infty)$ is such that
@@
Local well-posedness for (4.3) is provided by Theorem 1.3. The assumptions yield that the nonlinear potential energy density $F$ is well-approximated by the one of Ginzburg-Landau energy for $\rho$ close to $1$, see (2.16) and coercive. Further, there exists $10\leq\rho_{2}<$ 1 such that the negative part $F_{-}$ satisfies
@@
For (4.13), let $10<\rho_{1}<$ 1 denote the smaller root of $f$. Then, $10\leq\rho_{2}<\rho_{1}<$ 1.
@@
Moreover, there exists $C_{0}(\rho_{2})>0,C>0$ such that
@@
The first inequality immediately follows from Lemma 2.5. To show the second inequality, it suffices to prove that there exists $C_{2},C_{0}>0$ such that
@@
for all $|1+v|^{2}\geq R_{0}$. For $1+\delta\leq|1+v|^{2}\leq R_{0}$, it suffices to notice that $F$ is bounded from above and below away from 00 to conclude that there exists $C_{h}>0$ such that
@@
We observe that $\delta,\eta>0$ only depend on $10<\rho_{1}<$ 1 (and more precisely $\rho_{2}$) being the root of $f$ closest to but smaller than $1$. The expansion (2.16) which is determined by $\alpha_{1}$ and $g$ guaranties that $f$ has an isolated root in $1$. Hence, there exists $C_{0}=C_{0}(\eta)>0$ such that the claim follows. ∎
@@
In a first step, let $v_{0}\in\mathcal{F}_{1}$, i.e. $\psi_{0}:=1+v_{0}\in\mathbb{E}({\mathbf{R}}^{3})$ and $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$, such that $\Delta v_{0}\in L^{2}({\mathbf{R}}^{3})$, then $\psi=1+v\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{3}))$ and $\Delta v\in C([0,T];L^{2}({\mathbf{R}}^{3}))$ for all $0<T<T^{\ast}$ by virtue of Theorem 1.3. It follows
@@
upon integrating by parts and Young’s inequality. The second term is decomposed as
@@
with $\delta\in(0,1)$ such that (2.16) is valid for $||1+v|^{2}-1|\leq\delta$. We dispose of the terms separately. Note that on $\{|1+v|^{2}\leq 1-\delta\}$ one has $\operatorname{Re}(v)\in(-1-\sqrt{1-\delta},-1+\sqrt{1-\delta})$. Hence, for $\eta=1-\sqrt{1-\delta}$ we obtain
@@
where we used (2.17) in the last inequality. It remains to control $I_{3}$. In virtue of Assumption 4.7, it holds $F(\rho)>0$ for all $\rho>1$ and there exist $C>0$, $R_{0}>1$ such that $F(\rho)\geq C\rho^{1+\alpha_{1}}$ for all $\rho\geq R_{0}$. It follows,
@@
Under Assumption 4.7 and instead of $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$, one may alternatively assume that the initial data satisfies $\mathcal{H}(1+v_{0})$ and $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}^{2}$ sufficiently small adapting [42, Lemma 3.2] stated for cubic-quintic nonlinearities (1.18). Moreover, as pointed out in [43, Remark p. 2683] the same argument yields small data global well-posedness for the cubic-quintic nonlinearity where the quintic part is focusing and the cubic part defocusing, hence for $F$ being unbounded from below. Inspired, by this observation and the classical small data global well-posedness in $H^{1}$ for NLS eq, see e.g. [62, Chapter 3.4], we prove that (4.3) is globally well-posed in the energy space for small data provided that Assumption 1.5 holds.
@@
This proves Theorem 1.8. If $\operatorname{supp}(F_{-})\subset[0,1)$, then it suffices to assume $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}^{2}$ small instead of $\|\nabla v_{0}\|_{L^{2}}^{2}$ small.
@@
provided that $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}<<\eta^{\frac{3}{2}}$. Similarly, there exists $\nu>0$ only depending on $\delta>0$ as in (2.17) such that in $\operatorname{supp}(F\mathbf{1}_{\{|1+v|^{2}>1}\})$ it holds $|\operatorname{Re}(v_{0})|>\nu$ or $|\operatorname{Im}(v_{0})|>\nu$. Hence,
@@
Provided that $\varepsilon>0$ is sufficiently small, a continuity argument yields that $\mathcal{E}(1+v)(t)$ remains bounded, hence by virtue of the blow-up alternative states in Proposition 4.1 global existence follows. ∎

### v4

@@ 3.2. Global well-posedness
$\left\|\mathcal{N}_{2,\mathrm{int}}(\psi_{1})-\mathcal{N}_{2,\mathrm{int}}(\psi_{2})\right\|_{N^{0}}\leq\|\psi_{1}-\psi_{2}\|_{L_{t}^{1}L_{x}^{2}}+\|(|\psi_{1,\mathrm{int}}|^{2\alpha}+|\psi_{2,\mathrm{int}}|^{2\alpha})|c_{1}-c_{2}|\|_{L_{t}^{\frac{4}{3}}L_{x}^{\frac{3}{2}}}$
$+\|(|\psi_{1,\mathrm{int}}|^{2\alpha}+|\psi_{2,\mathrm{int}}|^{2\alpha})|v_{1}-v_{2}|\|_{L_{t}^{\frac{2}{3-\alpha}}L_{x}^{\frac{6}{2\alpha+1}}}$
@@
We now prove continuous dependence on the initial data. As in the proof of Proposition 3.2, we rely on an auxiliary metric to compensate for the lack of regularity of the nonlinearity $f$ and to deal with the non-integrability of the wave-functions. However, by virtue of Lemma 4.2, it suffices to consider the affine problem (4.3). This decomposition enables us to implement an argument in $L^{2}([0,T];L^{6}({\mathbf{R}}^{3}))$. In particular, it is sufficient to prove sequential continuity.
@@
where we used (2.24) in the second to last inequality and (2.20) to control the difference of the free solutions in the last inequality. More precisely,
@@
To bound the second contribution in (4.5), we proceed as in (3.21). More precisely, we observe that (3.24) remains valid upon replacing the admissible Strichartz pair $(4,4)$ for $d=2$ with $(\frac{8}{3},4)$ for $d=3$. Hence, the respective version of (3.24) reads that there exists $\theta_{2}\in(0,1]$ such that
@@
Summing up and applying the Strichartz estimate (2.24), we conclude from Lemma 4.4 that there exist $C=C(M)>0$ and $\theta>0$ such that
@@
as $n\rightarrow\infty$ and where $(q,r)$ as in (4.1). The proof follows closely the one of (3.17) to which we refer for full details. In view of the Strichartz estimates of Lemma 2.14 it follows that
@@
$\left|\nabla\mathcal{N}(\psi)\right|\leq C(1+|\psi|^{2\alpha})|\nabla\psi|\leq C(1+|\psi_{\mathrm{int}}|^{2\alpha})|\nabla\psi|.$
More precisely, for $G_{\mathrm{bd}},G_{\mathrm{int}}$ defined in (2.37) and upon applying (2.24), we split the non-homogeneous term:
@@
$\leq\|(G_{\mathrm{bd}})(\psi)|\nabla v-\nabla v_{n}|\|_{L_{t}^{1}L_{x}^{2}}+\|G_{\mathrm{int}}(\psi)|\nabla v-\nabla v_{n}|\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}$
$+\left\|\left(G_{\mathrm{bd}}(\psi)-G_{\mathrm{bd}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi)-G_{\mathrm{int}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}$
@@
$+\left\|\left(G_{\mathrm{bd}}(\psi)-G_{\mathrm{bd}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi)-G_{\mathrm{int}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}.$
@@
$\|\nabla v-\nabla v_{n}\|_{L^{\infty}([0,T_{2}],L^{2}({\mathbf{R}}^{3}))}+\|\nabla v-\nabla v_{n}\|_{L^{q}([0,T_{2}],L^{r}({\mathbf{R}}^{3}))}\leq C\delta(c+v_{0},c_{n}+v_{0}^{n})\\ +\left\|\left(G_{\mathrm{bd}}(\psi)-G_{\mathrm{bd}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi)-G_{\mathrm{int}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}.$
To conclude that (4.7) holds, it suffices to show that the second line of the right-hand side converges to 0 as $n$ goes to infinity. We proceed by contradiction assuming that there exists a subsequence still denoted $\psi_{n}$ such that there exists $\varepsilon>0$ such that for all $n$ sufficiently large,
(4.10) $\left\|\left(G_{\mathrm{bd}}(\psi)-G_{\mathrm{bd}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi)-G_{\mathrm{int}}(\psi_{n})\right)|\nabla v|\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\geq\varepsilon.$
Inequality (4.5) implies that up to extracting a further subsequence, still denoted $\psi_{n}$, that $\psi_{n}=c_{n}+v_{n}$ converges to $\psi=c+v$ a.e. on $0,T)\times{\mathbf{R}}^{3}$. By virtue of the Assumption 1.1, one has that $G_{\mathrm{bd}},G_{\mathrm{int}}$ are continuous. Therefore,
$\left|\left(G_{\mathrm{bd}}(\psi)-G_{\mathrm{bd}}(\psi_{n})\right)\right|\,|\nabla v|$ $\rightarrow 0\quad\text{a.e. in}\quad[0,T)\times{\mathbf{R}}^{3},$
$\left|\left(G_{\mathrm{int}}(\psi)-G_{\mathrm{int}}(\psi_{n})\right)\right|\,|\nabla v|$ $\rightarrow 0\quad\text{a.e. in}\quad[0,T)\times{\mathbf{R}}^{3}.$
@@
$\|G_{\mathrm{int}}(\psi_{n})\|_{L_{t}^{\infty}L_{x}^{\frac{2(\alpha+1)}{2\alpha}}({\mathbf{R}}^{3})}\leq C\||\psi_{n}|^{2\alpha}(1-\chi(\psi_{n}))\|_{L_{t}^{\infty}L_{x}^{\frac{2(\alpha+1)}{2\alpha}}({\mathbf{R}}^{3})}$
@@
for all $n\in{\mathbf{N}}$, where we exploited (2.4), namely that the measure of $\operatorname{supp}(1-\chi(\psi_{n}))$ is finite. We obtain that there exists $\phi\in L^{\infty}([0,T];L^{2(\alpha+1)}({\mathbf{R}}^{3}))$ such that $|\psi_{q,n}|\leq\phi$ a.e. on $0,T)\times{\mathbf{R}}^{3}$. Therefore, we control
$\left|\left(G_{\mathrm{bd}}(\psi)-G_{\mathrm{bd}}(\psi_{n})\right)\right|\,|\nabla v|\leq C|\nabla\psi|\in L^{1}([0,T);L^{2}({\mathbf{R}}^{3})),$
$\left|\left(G_{\mathrm{int}}(\psi)-G_{\mathrm{int}}(\psi_{n})\right)\right|\,|\nabla v|\leq C\left(|\psi|^{2\alpha}+|\phi|^{2\alpha}\right)|\nabla\psi|\in L^{q^{\prime}}([0,T);L^{r^{\prime}}({\mathbf{R}}^{3})).$
@@
Next, we show a persistence of regularity property and that the Hamiltonian energy $\mathcal{H}$ is conserved for regular solutions. The proof is completely analogous to the one for $d=2$, except that here we can exploit the affine structure of the energy space $\mathbb{E}$ and Sobolev embeddings which depend on the dimension. For the sake of clarity, we provide the proof of this lemma.
@@
By means of the Strichartz estimates of Lemma 2.14, for the admissible pair $(q,r)$ as in (4.1) and any $0<T<T^{\ast}$ we have that
@@
with $G_{1},G_{2}$ defined in (2.36). Upon splitting $G_{i}$ in $G_{i,\infty}$ and $G_{i,\mathrm{int}}$, as in (2.37), it follows that
@@
where we note that $0<\alpha-1<1$. It follows
@@
Finally, we conclude that $\mathcal{H}(c+v)(t)=\mathcal{H}(c+v_{0})$ by performing an analoguous argument as in the proof of Lemma 3.5 for $d=2$. ∎
@@
It only remains to show that the Hamiltonian energy is conserved for all solutions $\psi\in C([0,T^{\ast}),\mathbb{E}({\mathbf{R}}^{3}))$ which follows from Proposition 4.1 and an approximation by smooth solutions by means of Lemma 2.10 together with Lemma 4.5. ∎
@@ 4.1. Global well-posedness
Exploiting the affine structure of the energy space $\mathbb{E}({\mathbf{R}}^{3})$, we also prove global well-posedness for a class of equations for which the associated nonlinear potential energy density $F(|\psi|^{2})$ fails to be non-negative. Such equations arise for instance in nonlinear optics to investigate self-focusing phenomena in a defocusing medium, see [5, 49, 62]. A showcase model for such phenomena is (4.3) with competing subcritical power-type nonlinearities satisfying Assumption 1.5 and having the form
@@
where $a_{1},a_{2}>0$ and $0<\alpha_{2}<\alpha_{1}<2$. The defocusing nonlinearity is dominant for large intensities $|\psi|^{2}>>\rho_{0}$ and focusing phenomena occur for small intensities $|\psi|^{2}\leq\rho_{0}$ where $\rho_{0}$ is determined by the far-field. The case $\alpha_{1}=2$, $\alpha_{2}=1$ corresponds to the energy-critical cubic-quintic nonlinearity and is investigated in [45, 47]. As before, we set $\rho_{0}=1$, further as in (4.4), it suffices to consider $c=1$ and upon scaling space and time $a_{1}=1$. We are hence led to consider nonlinearities of the type
@@
where Assumption 1.5 implies $\frac{\alpha_{1}}{\alpha_{2}}>a_{2}$. Furthermore, we may assume that $a_{2}>1$ as otherwise $F\geq 0$. Indeed, the following hold
@@
    if $a_{2}>1$, then $f$ admits two positive real roots $\rho_{1},1$ with $0<\rho_{1}<1$ and $F$ displays a local minimum in $\rho=1$ and a local maximum in $\rho=\rho_{1}$. Depending on the location of the root $\rho_{1}$ two scenarios may occur:
@@
      the root $\rho_{1}$ is sufficiently close to 0 such that $F(\rho)\geq 0$ for all $\rho\geq 0$,
@@
Let $f$ be a real-valued function satisfying Assumption 1.5 and having the form
@@
where $0<\alpha_{1}<2$ and where $g\in C^{0}([0,\infty))\cap C^{1}(0,\infty)$ is such that
@@
Local well-posedness for (4.3) is provided by Theorem 1.3. The assumptions yield that the nonlinear potential energy density $F$ is well-approximated by the one of the Ginzburg-Landau energy for $\rho$ close to $1$, see (2.11) and coercive. Further, there exists $0\leq\rho_{2}<1$ such that the negative part $F_{-}$ satisfies
@@
For (4.13), let $0<\rho_{1}<1$ denote the smaller root of $f$. Then, $0\leq\rho_{2}<\rho_{1}<1$.
@@
Moreover, there exist $C_{0}(\rho_{1})>0$ and $C>0$ such that
@@
The first inequality immediately follows from Lemma 2.5. To show the second inequality, it suffices to prove that there exist $C_{2},C_{0}>0$ such that
@@
for all $|1+v|^{2}\geq R_{0}$. For $1+\delta\leq|1+v|^{2}\leq R_{0}$, it suffices to notice that $F$ is bounded from above and below away from 0 to conclude that there exists $C_{h}>0$ such that
@@
We observe that $\delta,\eta>0$ only depend on $0<\rho_{1}<1$ (and more precisely $\rho_{2}$) being the root of $f$ closest to but smaller than $1$. The expansion (2.11) which is determined by $\alpha_{1}$ and $g$ guaranties that $f$ has an isolated root in $1$. Hence, there exists $C_{0}=C_{0}(\eta)>0$ such that the claim follows. ∎
@@
First, let $v_{0}\in\mathcal{F}_{1}$, i.e. $\psi_{0}:=1+v_{0}\in\mathbb{E}({\mathbf{R}}^{3})$ and $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$, such that $\Delta v_{0}\in L^{2}({\mathbf{R}}^{3})$, then $\psi=1+v\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{3}))$ and $\Delta v\in C([0,T];L^{2}({\mathbf{R}}^{3}))$ for all $0<T<T^{\ast}$ by virtue of Theorem 1.3. It follows that
@@
upon integrating by parts and using Young’s inequality. The second term is decomposed as
@@
with $\delta\in(0,1)$ such that (2.11) is valid for $||1+v|^{2}-1|\leq\delta$. We treat of the terms separately. Note that on $\{|1+v|^{2}\leq 1-\delta\}$ one has $\operatorname{Re}(v)\in(-1-\sqrt{1-\delta},-1+\sqrt{1-\delta})$. Hence, for $\eta=1-\sqrt{1-\delta}$ we obtain
@@
where we used (2.12) in the last inequality. It remains to control $I_{3}$. By virtue of Assumption 4.7, we have that $F(\rho)>0$ for all $\rho>1$ and there exist $C>0$, $R_{0}>1$ such that $F(\rho)\geq C\rho^{1+\alpha_{1}}$ for all $\rho\geq R_{0}$. It follows that
@@
Under Assumption 4.7 and instead of $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$, one may alternatively assume that the initial data are such that $\mathcal{H}(1+v_{0})$ and $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}^{2}$ are sufficiently small adapting [45, Lemma 3.2] stated for cubic-quintic nonlinearities (1.18). Moreover, as pointed out in [46, Remark p. 2683] the same argument yields small data global well-posedness for the cubic-quintic nonlinearity where the quintic part is focusing and the cubic part defocusing, hence for $F$ being unbounded from below. Inspired, by this observation and the classical small data global well-posedness in $H^{1}$ for NLS eq, see e.g. [66, Chapter 3.4], we prove that (4.3) is globally well-posed in the energy space for small data provided that Assumption 1.5 holds.
@@
This proves Theorem 1.8. If $\operatorname{supp}(F_{-})\subset[0,1)$, it suffices to assume $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}^{2}$ small instead of $\|\nabla v_{0}\|_{L^{2}}^{2}$ small.
@@
provided that $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}<<\eta^{\frac{3}{2}}$. Similarly, there exists $\nu>0$ only depending on $\delta>0$ as in (2.12) such that in $\operatorname{supp}(F\mathbf{1}_{\{|1+v|^{2}>1}\})$, we have $|\operatorname{Re}(v_{0})|>\nu$ or $|\operatorname{Im}(v_{0})|>\nu$. Hence,
@@
Provided that $\varepsilon>0$ is sufficiently small, a continuity argument yields that $\mathcal{E}(1+v)(t)$ remains bounded, hence by virtue of the blow-up alternative stated in Proposition 4.1 global existence follows. ∎